3. The time taken to assemble a car in a certain plant has a normal distribution with mean of 25.4 hours and a standard deviation of 4.1 hours. Calculate the probability that a car can be assembled at this plant in the following period of time:


a) More than 28.8 hours


b) Between 18.6 and 27.5 hours


c) Between 25.0 and 34.0 hours


1
Expert's answer
2021-05-03T05:04:41-0400

"\mu = 25.4 \\

\sigma = 4.1"

a) P(X>28.8) = 1 -P(X<28.8)

"= 1 -P(Z< \frac{28.8 -25.4}{4.1}) \\

= 1 -P(Z<0.829) \\

= 1 -0.7964 \\

= 0.2036"

b) P(18.6<X<27.5) = P(X<27.5) -P(X<18.6)

"=P(Z< \frac{27.5-25.4}{4.1}) -P(Z< \frac{18.6-25.4}{4.1}) \\

= P(Z<0.512) -P(Z< -1.658) \\

= 0.6956 -0.0486 \\

= 0.647"

c) P(25.0<X<34.0) = P(X<34.0) -P(X<25.0)

"= P(Z< \frac{34.0-25.4}{4.1}) -P(Z< \frac{25.0 -25.4}{4.1}) \\

= P(Z<2.097) -P(Z<-0.097) \\

= 0.9820 -0.4613 \\

= 0.5207"